The specific heat capacity of ethanol is 2440J/kg °C. How many joules of energy will be required to heat 150g ethanol to 35°C if the starting temperature is 22°C?
1
Expert's answer
2019-03-26T06:19:49-0400

"E = cm{\Delta}t = 2440\frac{\text{J}}{\text{kg}^\circ\text{C}}\times0.15~\text{kg}\times(35-22)^\circ\text{C} = 4758~\text{J}. \\"